OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x^3y^4)2

jhonyy9 (jhonyy9):

hope you know that (2a^2)^2 = 2^2 *a^(2*2) = 4*a^4 1. when there is a product of ab on exponent x like (ab)^x so this mean that there is a on exponent x time b on exponent x when there is a on exponent x and all on exponent y like (a^x)^y so this mean that we need multiplie these exponents so we get a on exponent x time y so a^(x*y) hope that is clearly ,understandably

jhonyy9 (jhonyy9):

(3x^3y^4)^2 = 3^2 *x^3*2 *y^4*2 = 9*x^6*y^8 so the answer will be 9x^6y^8

jhonyy9 (jhonyy9):

hope that will help you